# Parallel Cluster-BFS and Applications to Shortest Paths

### v1

@@ Abstract
Breadth-first Search (BFS) is one of the most important graph processing subroutines, especially to compute the unweighted distance. Many applications may require to run BFS from multiple sources. Sequentially, when running BFS on a cluster of nearby vertices, a known optimization is to use bit-parallelism. Given a subset of vertices with size $k$ and the distance between any pair of them is no more than $d$, BFS can be applied to all of them in a total work of $O(dk/w+1)$, where $w$ is the length of a word in bits. We will refer to this approach as cluster-BFS (C-BFS). Such an approach has been studied and shown effective both in theory and in practice in the sequential setting. However, it remains unknown how this can be combined with thread-level parallelism for C-BFS.
@@ 1   Introduction
Many applications may require to run BFS from multiple sources. Examples of this are using BFS for low-diameter decomposition [37], all-pairs shortest paths (APSP), or oracles for exact or approximate APSP. A key observation is that bit-parallelism [20] can be used effectively when running BFS on a cluster of nearby vertices [20, 3]. Given a subset of vertices with size $k$ and the distance between any pair of them is no more than $d$, BFS can be applied to all of them in total $O(dk/w+1)$ work (number of operations), where $w$ is the length of a word in bits [20]. Since a machine word must hold at least $\Omega(\log n)$ bits to store a pointer, this means $w=\Omega(\log n)$. We will refer to this approach as cluster-BFS (C-BFS), and present more details in Sec. 3. In 2012, Chan [20] first observed this and used this idea to develop an all-pair shortest-path algorithm that runs in $O(mn/w)$ work (when $m=\Omega(n\log n\log\log\log n))$. In addition to saving time, C-BFS also saves space: it only uses $O(d)$ words per $w$ vertices instead of a word (or at least enough bits to store a distance) per vertex as in standard BFS. In 2013, Akiba et al. [3] used this idea in the exact two-hop distance oracle but only considered the special case for $d=2$ (a star-shaped cluster: a vertex and its neighbors). We refer to this algorithm as the AIY algorithm. Both of the previous papers focus on the sequential setting.
While the C-BFS with bit-level parallelism has shown effective in sequential settings, surprisingly, we know of no previous work studying the (thread-level) parallelism on it. BFS is one of the most well-studied parallel graph processing problems, and state-of-the-art solutions have been highly optimized using techniques such as direction optimizations [8, 45]. To be practical, any C-BFS would have to compete with these. Our goal is to develop efficient C-BFS with high parallelism such that it 1) achieves the same level of parallelism as the standard parallel BFS, with additional benefits by using clusters, 2) supports a clean interface that is flexible for different parameter settings (i.e., varying $d$ and $k$), and 3) facilitates various real-world applications. In this paper, we provide a systematic study of parallel C-BFS and achieve all three goals above.
@@
To achieve a flexible interface, we designed our algorithm to work for general $k$ and $d$ and can be easily integrated into different applications using user-defined parameters.
@@
We implemented our C-BFS algorithm and the two applications. We compared our algorithm with multiple baselines to study the performance gain in depth. We tested 18 graphs with various types and sizes on a 96-core machine. Compared to standard sequential BFS, our algorithm employs both bit-parallelism (on clusters) and thread-level parallelism (along with optimizations used in parallel BFS) to improve performance. In the simplest setting where $k=64$ and $d=2$, the combination of them enables up to 1119$\times$ speedup (500$\times$ on average) compared to the plain sequential BFS, where bit-level and thread-level parallelism each contributes about 20$\times$ speedup. Interestingly, by comparing with the performance of existing work (Ligra [45], where only thread-parallelism is used, and AIY [4], where only bit-level parallelism is used), we observed that both bit-level parallelism and thread-level parallelism and work very well in synergy. Each of them still fully contributes to the performance when the other presents, achieving the same level of improvement as when used independently (see Fig. 1). Therefore, we believe our work on an efficient implementation combining thread-level and bit-level parallelism fills the gap in the existing study of both C-BFS and parallel BFS.
@@ 2   Preliminaries
Computational Model.   We use the binary fork-join parallelism [23, 10], with work-span analysis [14, 28]. We assume a set of threads that access a shared memory. A process can fork two child software threads to work in parallel. When both children complete, the parent process continues. A parallel for-loop can be simulated by recursive forks in logarithmic levels. The work of an algorithm is the total number of instructions, and the span is the length of the longest sequence of dependent instructions. We can execute the computation using a randomized work-stealing scheduler [14, 27] in practice.
@@
Parallel BFS.   We briefly review parallel BFS, because it is one of our baselines and some of the concepts are also used in our cluster-BFS. Parallel BFS starts from a single source $s\in V$ (high-level idea in Alg. 3). The algorithm maintains a frontier of vertices to explore in each round, starting from the source, and finish in $D$ rounds. In round $i$, the algorithm processes (visits their neighbors) of the current frontier $\mathcal{F}_{i}$, and puts all their (unvisited) neighbors to the next frontier $\mathcal{F}_{i+1}$. If multiple vertices in $\mathcal{F}_{i}$ attempt to add the same vertex to $\mathcal{F}_{i+1}$, a compare_and_swap is used to guarantee that only one will succeed.
One widely-used optimization for parallel BFS is the directional optimization [8, 45]. At a high level, when the frontier size $|\mathcal{F}_{i}|$ is large, the algorithm will not process $\mathcal{F}_{i}$, but instead visit each unprocessed vertex $v$, and determine if $v$ has an incoming neighbor in $\mathcal{F}_{i}$. If so, $v$ will be put in $\mathcal{F}_{i+1}$. Such an optimization is observed to be effective, especially on small-diameter graphs. We present more details in Sec. A.
@@ 3   Parallel Cluster-BFS
Cluster-BFS (C-BFS) runs BFS from a cluster of sources $S\subseteq V$. If the sources have diameter $d$ (maximum distance between any pair), then all distances from $S$ to any vertex $v\in V$ will differ by at most $d$. This means that if we run a set of BFSs from $S$, synchronously, every $v\in V$ will appear in at most $d+1$ consecutive frontiers. Cluster-BFS takes advantage of this by representing all the sources in $S$ that can visit a given vertex, on a given round, as a vector of booleans (bits). In this way, each vertex will be visited at most $d+1$ times instead of $|S|$ times if all searches are performed separately. We will explain more details in section 3.1. Importantly, if the bit-vector fits in $O(1)$ words, the distances of all $|S|$ sources can be handled (propagated from a vertex in the current frontier to a neighbor) with $O(1)$ bitwise logical operations. Since a machine word must hold at least $\Omega(\log n)$ bits (so it can represent a pointer), this means the algorithm can save a factor of $\Omega(\log n/d)$ work.
It appears that Chan first described this idea [20], but he did not go into any details of the implementation, but just saying that this is possible. Akiba et al. [3] later showed a concrete implementation based on this idea, with the limitations that it is sequential and only works on the cluster of a star with $d=2$ (a center vertex and its neighbors). To the best of our knowledge, there has been no previous work on developing a parallel implementation of the cluster-BFS algorithm. Indeed this is pretty challenging to achieve high performance given how BFS has been widely studied with numerous optimizations both sequentially and in parallel. In this paper, we propose our algorithm, given as Alg. 1, which is an efficient parallel cluster-BFS implementation with a general interface and low coding effort. In the following sections, we will first introduce the bitwise representation to maintain the distances in cluster-BFS and then give our parallel cluster-BFS algorithm.
@@ 3.1   Cluster Distance Representations
Note that if $|S|=w$, we can use one word of bit-vector to represent any subset of $S^{\prime}\subseteq S$: bit $i$ is 1 iff. the $i$-th element in $S$ is also in $S^{\prime}$. We call such a bit-flag representation of a subset of $S$ a bit-subset. In this way, a cluster distance vector only takes $d+1$ words for bit-subsets and one byte for distance to store the shortest distances from $v$ to $|S|$ sources.
@@
The main idea of cluster-BFS is to use bitwise operations on these bit-subsets to quickly compute the union/intersection of the sets, so that we can use the cluster distance vector of $v$ to compute the cluster distance vectors of its neighbor $u$ in constant time. In the following, we elaborate on our parallel cluster-BFS algorithm.
@@ 3.2   Our Parallel Algorithm
Our cluster-BFS algorithm is based on the following fact: if $u$ and $v$ are neighbors and there is a path from a source $s\in S$ to $u$ with length $i-1$, then there must exist a path from $s$ to $v$ with length $i$. In round $i$, $u$ records all the sources in $S$ that reach it in round $i$ by a bit-subset. When $u$ visits its neighbor $v$, $u$ propagates this bit-subset to $v$ by taking a bitwise Or operation with the bit-subset representing the vertices reaching $v$ in round $i+1$ (Alg. 1: Alg. 1). According to Cor. 3.1, all the vertices in $S$ will visit $v$ at least once during round $\Delta_{v}$ to round $\Delta_{v}+d$; in other words, all the vertices will be put into the frontier for $d+1$ times. Therefore, we need to record the number of times $v$ has been put in the frontier. When $v$ has been put in the frontier for $d+1$ times, since all sources in $S$ must have already visited $v$, we do not need to process $v$ anymore. Otherwise, we will process $v$ and put it to the next frontier since other sources in $S$ may visit $v$ in the future.
Our cluster-BFS algorithm is given in Alg. 1. In our algorithm, we use a boolean array $S_{seen}[\cdot]$ to store whether a vertex has been visited in all previous rounds by any sources in $S$, and $S_{next}[v]$ includes the vertex if it is also in the current frontier (i.e., visited by any vertex in the current round). We denote $\Delta_{v}$ as the first round that any vertex from $S$ touches vertex $v$. Alg. 1 has two stages: initialization and traversing.
@@
Traversing   At the beginning, the algorithm puts all the sources $s\in S$ into the first frontier $\mathcal{F}_{0}$. Then, we visit all vertices by frontiers. In each round, we process frontiers in two stages, where the first stage processes vertices and the second stage processes edges. In the first stage (Alg. 1 to Alg. 1), we first compute the sources that newly visited $u$ by $S_{new}\leftarrow S_{next}[u]\setminus S_{seen}[u]$ (Alg. 1). Note that the bit-subset $S_{new}$ contains sources whose distances are $i$, which is also ${S}_{u}[i-\Delta_{u}]$ (Alg. 1). Then we update $S_{seen}[u]$ to include newly visited vertices (Alg. 1), and set $\Delta_{u}$ to the current round number if it has not been set yet (Alg. 1). In the second stage (Alg. 1 to Alg. 1), we process the neighbor vertices of the current frontier that have not been visited for $d$ times already (Alg. 1). For an edge from $u\in\mathcal{F}_{i}$ to its neighbor $v$, we propagate the sources seen so far by $u$, $S_{seen}[u]$, to $v$ (Alg. 1). In general, if any source $s\in S$ visited $u$ in the previous round, $s$ should also visit $v$ in this round, and should be included in the $S_{next}[v]$ for $v$ in this round. If the $S_{next}[v]$ is changed, which means there are new sources visiting $v$, $v$ should be added to the next frontier. To avoid duplication in the next frontier, only the one that can successfully set $r[v]$ to $i$ (Alg. 1) by compare_and_swap will put $v$ to the next frontier. Note that we can further benefit from the directional optimization that is commonly used in parallel BFS. Additional details about the directional optimization are available in Sec. A.
@@
Given the bit-subset subsets $S_{1},S_{2}$ of a set $S$ with size $k$, we can compute the bit-subset representation of $S_{1}\cup S_{2}$, $S_{1}-S_{2}$ in $O(k/w+1)$ work and $O(\log(k/w)+1)$ span, where $w$ is the word length.
@@
    A bit-subset with $k$ bits needs $\lceil k/w\rceil$ words. These $\lceil k/w\rceil$ words can be processed in parallel. With the constant cost for bitwise or/not operations for a single word, the work and span for $S_{1}\cup S_{2}$, $S_{1}-S_{2}$ are $O(k/w+1)$ and $O(\log(k/w)+1)$.
@@
If we take $k=O(w)$, such that each bit-subset fits within a constant number of words, the work simplifies to $O(dm)$ and the span remains $O((D+d)\log n)$, which matches the work and span of a single BFS. Since $w=O(\log n)$, this means that we can compute $O(\log n)$ more BFSs with asymptotically the same cost.
@@ 4   Applications
In this section, we show two applications that can benefit from our new parallel Cluster-BFS algorithm. Both applications are distance oracles (DO) for unweighted graphs. A distance oracle (DO) is an index designed to answer the shortest distances between two vertices on a graph. Although such a query can always be answered by computing the distance on the fly (e.g., running a BFS from one of the query vertices), it can be inefficient for applications requiring low latency or requesting multiple queries. A distance oracle aims to store certain information by preprocessing to accelerate the distance queries.

### v2

@@ Abstract
Breadth-first Search (BFS) is one of the most important graph processing subroutines, especially for computing the unweighted distance. Many applications may require running BFS from multiple sources. Sequentially, when running BFS on a cluster of nearby vertices, a known optimization is using bit-parallelism. Given a subset of vertices with size $k$ and the distance between any pair of them is no more than $d$, BFS can be applied to all of them in total work $O(dm(k/w+1))$, where $w$ is the length of a word in bits and $m$ is the number of edges. We will refer to this approach as cluster-BFS (C-BFS). Such an approach has been studied and shown effective both in theory and in practice in the sequential setting. However, it remains unknown how this can be combined with thread-level parallelism.
@@ 1   Introduction
Many applications may require running BFS from multiple sources. Examples of this are using BFS for low-diameter decomposition [37], all-pairs shortest paths (APSP), or oracles for exact or approximate APSP. A key observation is that bit-parallelism [20] can be used effectively when running BFS on a cluster of nearby vertices [20, 3]. Given a subset of vertices with size $k$ and the distance between any pair of them is no more than $d$, BFS can be applied to all of them in total $O(dm(k/w+1))$ work (number of operations), where $w$ is the length of a word in bits and $m$ is the number of edges [20]. Since a machine word must hold at least $\Omega(\log n)$ bits to store a pointer, this means $w=\Omega(\log n)$. We will refer to this approach as cluster-BFS (C-BFS), and present more details in Sec. 3. Chan [20] used this idea to develop an all-pair shortest-path algorithm that runs in $O(mn/w)$ work (when $m=\Omega(n\log n\log\log\log n))$. In addition to saving time, C-BFS also saves space: it only uses $O(d)$ words per $w$ vertices instead of a word (or at least enough bits to store a distance) per vertex as in standard BFS. Akiba et al. [3] used this idea in the exact two-hop distance oracle but only considered the special case for $d=2$ (a star-shaped cluster: a vertex and its neighbors). We refer to this algorithm as the AIY algorithm. Both of the previous papers focus on the sequential setting.
While the C-BFS with bit-level parallelism has shown to be effective in sequential settings, surprisingly, we know of no previous work combining it with thread-level parallelism. BFS is one of the most well-studied parallel graph processing problems, and state-of-the-art solutions have been highly optimized using techniques such as direction optimizations [8, 45]. To be practical, any C-BFS would have to compete with these. Our goal is to develop an efficient C-BFS with high parallelism such that it (1) achieves the same level of parallelism as the standard parallel BFS, with additional benefits by using clusters, (2) supports a clean interface that is flexible for different parameter settings (i.e., varying $d$ and $k$), and (3) facilitates various real-world applications. In this paper, we provide a systematic study of parallel C-BFS and achieve all three goals above.
@@
To achieve a flexible interface, we designed our algorithm for general $k$ and $d$, easily integrating into various applications with user-defined parameters.
@@
We implemented our C-BFS algorithm and the two applications. We compare our algorithm with multiple baselines to study the performance gain in depth, and test it on 18 graphs with various types and sizes on a 96-core machine. Compared to standard sequential BFS, our algorithm employs both bit-parallelism (on clusters) and thread-level parallelism (along with optimizations used in parallel BFS) to improve performance. In the simplest setting where $k=64$ and $d=2$, the combination of them enables up to 1119$\times$ speedup (500$\times$ on average) compared to the plain sequential BFS, where bit-level and thread-level parallelism each contributes about 20$\times$ speedup. Interestingly, by comparing with the performance of existing work (Ligra [45], where only thread-parallelism is used, and AIY [4], where only bit-level parallelism is used), we observed that both bit-level parallelism and thread-level parallelism and work very well in synergy. Each of them still fully contributes to the performance when the other is present, achieving the same level of improvement as when used independently (see Fig. 1). Therefore, we believe our work on an efficient implementation combining thread-level and bit-level parallelism fills the gap in the existing study of both C-BFS and parallel BFS.
@@ 2   Preliminaries
Computational Model.   We use the binary fork-join parallel model [23, 10], with work-span analysis [14, 28]. We assume a set of threads that access a shared memory. A thread can fork two child threads to work in parallel, and then waits. When both children complete, the parent thread continues. A parallel for-loop can be simulated by recursive forks in logarithmic levels. The work of an algorithm is the total number of instructions, and the span is the length of the longest sequence of dependent instructions. We can execute the computation using a randomized work-stealing scheduler [14, 27].
@@
Parallel BFS.   We briefly review parallel BFS, because it is one of our baselines, and some of the concepts are also used in our cluster-BFS. Parallel BFS starts from a single source $s\in V$ (high-level idea in Alg. 3). The algorithm maintains a frontier of vertices to explore in each round, starting from the source, and finishes in at most $D$ rounds. In round $i$, the algorithm processes (visits their neighbors) of the current frontier $\mathcal{F}_{i}$, and puts all their (unvisited) neighbors in the next frontier $\mathcal{F}_{i+1}$. If multiple vertices in $\mathcal{F}_{i}$ attempt to add the same vertex to $\mathcal{F}_{i+1}$, a compare_and_swap is used to guarantee that only one will succeed.
One widely-used optimization for parallel BFS is directional optimization [8, 45]. At a high level, when the frontier size $|\mathcal{F}_{i}|$ is large, the algorithm will not process $\mathcal{F}_{i}$, but instead visit each unprocessed vertex $v$, and determine if $v$ has an incoming neighbor in $\mathcal{F}_{i}$. If so, $v$ will be put in $\mathcal{F}_{i+1}$. Such an optimization is observed to be effective, especially on small-diameter graphs. We present more details in Sec. A.
@@ 3   Parallel Cluster-BFS
Cluster-BFS (C-BFS) runs BFS from a cluster of sources $S\subseteq V$. If the sources have diameter $d$ (maximum distance between any pair), then all distances from $S$ to any vertex $v\in V$ will differ by at most $d$. This means that if we run a set of BFSs from $S$, synchronously, every $v\in V$ will appear in at most $d+1$ consecutive frontiers. Cluster-BFS takes advantage of this by representing all the sources in $S$ that can visit a given vertex, on a given round, as a vector of booleans (bits). In this way, each vertex will be visited at most $d+1$ times instead of $|S|$ times if all searches are performed separately. More details are described in section 3.1. Importantly, if the bit-vector fits in $O(1)$ words, the distances of all $|S|$ sources can be handled (propagated from a vertex in the current frontier to a neighbor) with $O(1)$ bitwise logical operations. Since a machine word must hold at least $\Omega(\log n)$ bits (so it can represent a pointer), this means the algorithm can save a factor of $\Omega(\log n/d)$ work.
It appears that Chan first described this idea [20], but he did not go into any details of the implementation, but just saying that this is possible. Akiba et al. [3] later showed a concrete implementation based on this idea, with the limitations that it is sequential and only works on the cluster of a star with $d=2$ (a center vertex and its neighbors). To the best of our knowledge, there has been no previous work on developing a parallel implementation of the cluster-BFS algorithm. Indeed it is challenging to achieve high performance given how BFS has been widely studied with numerous optimizations both sequentially and in parallel. In this paper, we propose our algorithm, given as Alg. 1, which is an efficient parallel cluster-BFS implementation with a general interface and low coding effort. In the following sections, we will first introduce the bitwise representation to maintain the distances in cluster-BFS and then give our parallel cluster-BFS algorithm.
@@ 3.1   Cluster Distance Representations
Note that if $|S|=w$, we can use a one word bit-vector to represent any subset of $S^{\prime}\subseteq S$: bit $i$ is 1 iff. the $i$-th element in $S$ is also in $S^{\prime}$. We call such a representation of a subset of $S$ a bit-subset. In this way, a cluster distance vector only takes $d+1$ words for bit-subsets and one byte to store the shortest distances from $v$ to $|S|$ sources (assuming $D<256$).
@@
The main idea of cluster-BFS is to use bitwise operations on bit-subsets to quickly compute the union/intersection of the sets, allowing us to use the cluster distance vector of $v$ to compute the cluster distance vectors of its neighbor $u$ in constant time. In the following, we elaborate on our parallel cluster-BFS algorithm.
@@ 3.2   Our Parallel Algorithm
Our cluster-BFS algorithm is based on the following fact: if $u$ and $v$ are neighbors and there is a path from a source $s\in S$ to $u$ with length $i-1$, then there must exist a path from $s$ to $v$ with length $i$. In round $i$, $u$ records all the sources in $S$ that reach it in round $i$ by a bit-subset. When $u$ visits its neighbor $v$, $u$ propagates this bit-subset to $v$ by taking a bitwise Or operation with the bit-subset representing the vertices reaching $v$ in round $i+1$ (Alg. 1: Alg. 1). According to Cor. 3.1, all the vertices in $S$ will visit $v$ at least once during round $\Delta_{v}$ to round $\Delta_{v}+d$; in other words, all the vertices will be put into the frontier for $d+1$ times. Therefore, we need to record the number of times $v$ has been put in the frontier. When $v$ has been put in the frontier $d+1$ times, since all sources in $S$ must have already visited $v$, we do not need to process $v$ anymore. Otherwise, we will process $v$ and put it to the next frontier since other sources in $S$ may visit $v$ in the future.
In our algorithm, we use a boolean array $S_{seen}[\cdot]$ to store whether a vertex has been visited in all previous rounds by any sources in $S$, and $S_{next}[v]$ includes the vertex if it is also in the current frontier (i.e., visited by any vertex in the current round). We denote $\Delta_{v}$ as the first round that any vertex from $S$ touches vertex $v$. Alg. 1 has two stages: initialization and traversing.
@@
Traversing   At the beginning, the algorithm puts all the sources $s\in S$ into the first frontier $\mathcal{F}_{0}$. Then, we visit all vertices by frontiers. In each round, we process frontiers in two stages, where the first stage processes vertices and the second stage processes edges. In the first stage (Alg. 1 to Alg. 1), we first compute the sources that newly visited $u$ by $S_{new}\leftarrow S_{next}[u]\setminus S_{seen}[u]$ (Alg. 1). Note that the bit-subset $S_{new}$ contains sources whose distances are $i$, which is also ${S}_{u}[i-\Delta_{u}]$ (Alg. 1). Then we update $S_{seen}[u]$ to include newly visited vertices (Alg. 1), and set $\Delta_{u}$ to the current round number if it has not been set yet (Alg. 1). In the second stage (Alg. 1 to Alg. 1), we process the neighbor vertices of the current frontier that have not been visited for $d$ times already (Alg. 1). For an edge from $u\in\mathcal{F}_{i}$ to its neighbor $v$, we propagate the sources seen so far by $u$, $S_{seen}[u]$, to $v$ (Alg. 1). In general, if any source $s\in S$ visited $u$ in the previous round, $s$ should also visit $v$ in this round, and should be included in the $S_{next}[v]$ for $v$ in this round. If the $S_{next}[v]$ is changed, which means there are new sources visiting $v$, $v$ should be added to the next frontier. To avoid duplication in the next frontier, only the one that can successfully set $r[v]$ to $i$ (Alg. 1) by compare_and_swap will put $v$ to the next frontier. Note that we can further benefit from the directional optimization that is commonly used in parallel BFS. Additional details about the directional optimization are given in Sec. A.
@@
Given the bit-subset subsets $S_{1}$ and $S_{2}$ of a set $S$ with size $k$, we can compute the bit-subset representation of $S_{1}\cup S_{2}$, $S_{1}\setminus S_{2}$ in $O(k/w+1)$ work and $O(\log(k/w)+1)$ span, where $w$ is the word length.
@@
    A bit-subset with $k$ bits needs $\lceil k/w\rceil$ words. These $\lceil k/w\rceil$ words can be processed in parallel. With the constant cost for bitwise or/not operations for a single word, the work and span for $S_{1}\cup S_{2}$, $S_{1}\setminus S_{2}$ are $O(k/w+1)$ and $O(\log(k/w)+1)$.
@@
If we take $k=\Theta(w)$, such that each bit-subset fits within a constant number of words, the work simplifies to $O(dm)$ and the span remains $O((D+d)\log n)$, which matches the work and span of a single BFS. Since $w=\Omega(\log n)$, this means that we can compute $O(\log n)$ more BFSs with asymptotically the same cost.
@@ 4   Applications
In this section, we show two applications that can benefit from our new parallel Cluster-BFS algorithm. Both applications are distance oracles (DO) for unweighted graphs. A distance oracle (DO) is an index designed to answer the shortest distances between two vertices on a graph. Although such a query can always be answered by computing the distance on the fly (e.g., running a BFS from one of the query vertices), this can be inefficient for applications requiring low latency or requesting multiple queries. A distance oracle aims to store information generated during preprocessing to accelerate the distance queries.